# Surrogate-based multilevel Monte Carlo methods for uncertainty quantification in the Grad-Shafranov free boundary problem

### v1

@@ 1.1 The Grad-Shafranov free boundary problem with uncertainty
$-\nabla\,\cdot\,\left(\frac{1}{\mu(u,\boldsymbol{\omega})r}\nabla u(\cdot,\boldsymbol{\omega})\right)=\left\{\begin{array}[]{ll}\frac{d}{du}p(u,\boldsymbol{\omega})+\frac{1}{2\,\mu r}\frac{d}{du}g^{2}(u,\boldsymbol{\omega})&\quad\text{ in }\Omega_{p}(u,\boldsymbol{\omega})\,,\\ I_{k}(\boldsymbol{\omega})/S_{k}&\quad\text{ in }\Omega_{W_{k}}\\ 0&\quad\text{ elsewhere.}\end{array}\right.$ (2)
@@ 1.2 Overview of the article
In its simplest form, a Monte Carlo estimate of the expectation $\mathbb{E}(u)$ gathers a large number, $N$, of the realizations of $u(\cdot,\boldsymbol{\omega})$ and estimates the expectation by the mean of the sample as
$\mathbb{E}(u)\approx\frac{1}{N}\sum_{i=1}^{N}u\left(\cdot,\boldsymbol{\omega}^{(i)}\right)\,.$
Based on the sample mean, the approximation is unbiased and, by the central limit theorem, it will converge to the true value as $N^{-1/2}$. Due to the slow rate of convergence, a large number of samples may need to be gathered to obtain a reasonably accurate approximation of $\mathbb{E}(u)$. Therefore, in simple words, the cost of obtaining a Monte Carlo estimate boils down to the workload associated with sampling. In our case, $u$ maps the stochastic parameters in (2) to the solution of the free boundary problem. In practical terms, the task of obtaining a Monte Carlo approximation for $\mathbb{E}(u)$ entails the use of a computational discretization of the mapping, and the computation of a large number of numerical solutions to (2) for different values of the stochastic parameters. Unfortunately, even if obtaining such a pool of numerical solutions is computationally feasible, the cost of numerically solving the free boundary problem repeatedly can make the process unpractical. The goal then is then to bypass the need for a direct numerical solution of (2) for every new sample gathered, or at least mitigate the work required to obtain each sample.
@@ 2.1 Sparse grid stochastic collocation
where $P$ denotes the number of nodes in the sparse grid, $C$ is a positive constant and the power $\nu$ is an increasing function of the size of the domain of definition of the function’s analytic extension in the complex plane. On the other hand, if the mapping is less regular with respect to the stochastic parameter $\boldsymbol{\omega}$, for instance of class $C^{k}(\boldsymbol{W},Z)$, then [2, Theorem 8]
@@ 2.2 Notational considerations
$\mathbb{V}(u):=\mathbb{E}\left(\left\|u-\mathbb{E}(u)\right\|_{Z}^{2}\right)\,,$ (8)
@@
where $\alpha$ and $\nu$ are the orders of the sample-wise discretization error and interpolation error in (6); the constants $C_{m}(\boldsymbol{\omega}^{(i)})$ and $C_{p}(\boldsymbol{\omega}^{(i)})$ depend only on the problem geometry and the particular realization $\boldsymbol{\omega}^{(i)}$. We remark that in what follows, when describing the different approaches to surrogate building, the associated values of the constants $C_{m}$, $C_{p}$, $\alpha$, and $\nu$ may vary from section to section.
@@ 2.3 Collocation with a single level of spatial discretization
Considering that, for a nonlinear problem like the one in our application, the computational cost of obtaining a sample is itself a random variable, we will estimate the computational work required to build this kind of surrogate, by assuming that the average work per sample is of the form $W=\mathcal{O}(M^{\gamma})$ for some $\gamma>0$ that depends on the particular method used to solve the system of equations arising from the discretization. Hence, making use of (12), for a surrogate built using $P$ sparse grid nodes, the expected work to construct the surrogate, measured in terms of $\epsilon$, is given by
@@
where the subscript “SL-SL” denotes the surrogate built with a single-level sparse grid and a single-level spatial grid, the superscript “off” represents an offline process, and, to recall, $\epsilon$ represents the desired accuracy, $\alpha$ is the convergence rate of the spatial discretization, $\nu$ is the convergence rate of the interpolant, $\gamma$ is the rate associated with the non-linear solver that characterizes the average work per sample, and $A\simeq B$ means $C_{1}B\leq A\leq C_{2}B$ for positive $A$ and $B$, with constants $C_{1},C_{2}$ independent of sample size $N$ and number of spatial grid nodes $M$.
@@ 2.4 Collocation with multilevel spatial discretization
$\left\|Y_{\ell}^{(i)}-\widehat{Y}_{\ell}^{(i)}\right\|_{Z}=\left\|\left(u_{\ell}^{(i)}-u_{\ell-1}^{(i)}\right)-\left({{\color[rgb]{0,0,0}\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}}}\right)\right\|_{Z}=\left\|\left(u_{\ell}^{(i)}-\widehat{u}_{\ell}^{(i)}\right)-\left(u_{\ell-1}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)\right\|_{Z}$
@@
These intuitive observations in fact hold for the case of Lagrangian interpolation and Smolyak grids based on nested quadrature points—which include the Clenshaw-Curtis strategy we employ in this paper. Moreover, it was shown in [35, Section 5] that there are in fact positive constants $C^{(i)},\rho$, and $\nu$ such that
@@ 3.1 Sampling with a single-spatial level surrogate
$A\left(\widehat{u}_{h}\right):=\frac{1}{N}\sum_{i=1}^{N}\widehat{u}_{h}^{(i)}.$
@@
$\mathbb{E}(A(\widehat{u}_{h}))=\mathbb{E}(\widehat{u}_{h})\qquad\text{ and }\qquad\mathbb{V}(A(\widehat{u}_{h}))=\mathbb{V}(\widehat{u}_{h})/N\,,$ (23)
where $\mathbb{V}(\cdot)$ is defined as in (8). To assess the accuracy of the estimator we will use the mean squared error (MSE) defined as
$\mathcal{E}_{A}^{2}:=\mathbb{E}\left[\left\|\mathbb{E}(u)-A\left(\widehat{u}_{h}\right)\right\|_{Z}^{2}\right]\,$
@@
$\mathcal{E}_{A}^{2}=\mathbb{E}\left[\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|^{2}_{Z}\right]+\mathbb{E}\left[\left\|\mathbb{E}\left(\widehat{u}_{h}\right)-A\left(\widehat{u}_{h}\right)\right\|_{Z}^{2}\right]=\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|^{2}_{Z}+\frac{\mathbb{V}\left(\widehat{u}_{h}\right)}{N}=\mathcal{E}_{\text{Bias}}^{2}+\mathcal{E}_{\text{Stat}}^{2},$
@@
$\mathcal{E}_{\text{Bias}}^{2}:=\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|^{2}_{Z}\qquad\text{ and }\qquad\mathcal{E}_{\text{Stat}}^{2}:=\frac{\mathbb{V}\left(\widehat{u}_{h}\right)}{N}.$
The bias term can be further bounded by the sum of discretization error, $\mathcal{E}_{\text{Dis}}:=\|\mathbb{E}(u)-\mathbb{E}(u_{h})\|_{Z}$, and interpolation error  $\mathcal{E}_{\text{Interp}}:=\left\|\mathbb{E}(u_{h})-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|_{Z}$, as
$\mathcal{E}_{\text{Bias}}=\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|_{Z}\leq\|\mathbb{E}(u)-\mathbb{E}(u_{h})\|_{Z}+\left\|\mathbb{E}(u_{h})-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|_{Z}=\mathcal{E}_{\text{Dis}}+\mathcal{E}_{\text{Interp}}\,.$
@@
If the expectation of $u$ is nonzero, the normalized mean squared error (nMSE), denoted ¯$\overline{\mathcal{E}}_{A}^{2}$, is considered. A similar normalization and notation is used for the discretization, interpolation and statistical errors. When the normalizing factor $\left\|\mathbb{E}(u)\right\|_{Z}^{2}$ is not available or computable we approximate it by $\left\|\mathbb{E}(u_{h})\right\|_{Z}^{2}$. For instance the nMSE is approximated as
¯$\overline{\mathcal{E}}_{A}^{2}\approx\frac{\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{h}\right)\right\|^{2}_{Z}}{\left\|\mathbb{E}(u_{h})\right\|_{Z}^{2}}+\frac{\mathbb{V}\left(\widehat{u}_{h}\right)}{N\left\|\mathbb{E}(u_{h})\right\|_{Z}^{2}}\leq\left(\overline{\mathcal{E}}_{\text{Dis}}+\overline{\mathcal{E}}_{\text{Interp}}\right)^{2}+\overline{\mathcal{E}}_{\text{Stat}}^{2}.$
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\theta_{2}\epsilon,\qquad\text{ and }\qquad\overline{\mathcal{E}}_{\text{Stat}}=\sqrt{\frac{\mathbb{V}\left(\widehat{u}_{h}\right)}{N\left\|\mathbb{E}(u_{h})\right\|_{Z}^{2}}}\leq\sqrt{\theta}\epsilon,$ (24)
where $\theta\in(0,1)$ is the splitting parameter between the relative bias and the statistical errors, and $\theta_{1}$ and $\theta_{2}$ are the splitting parameters between the discretization and interpolation errors, such that $\theta_{2}=\sqrt{1-\theta}-\theta_{1}$. The number of grid points in the physical space $M$ and the parameter space $P$ are determined by (12) with a scaling factor $\left\|\mathbb{E}(u_{h})\right\|_{Z}$, while the sample size $N$ is estimated from (24) as
$N=\left\lceil\frac{\mathbb{V}\left(\widehat{u}_{h}\right)}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{h})\right\|_{Z}^{2}}\right\rceil\simeq\epsilon^{-2}.$ (25)
@@ 3.2 Sampling with surrogates with multilevel spatial discretization
We express the surrogate at the finest spatial grid level $L$ as a telescoping sum of surrogates on coarser grids, and then exploit the linearity of the expectation to approximate $\mathbb{E}(u_{h})$ as
$\mathbb{E}(u_{h})\approx\mathbb{E}\left(\widehat{u}_{L}\right)=\sum_{\ell=0}^{L}\mathbb{E}(\widehat{u}_{\ell}-\widehat{u}_{\ell-1})=\sum_{\ell=0}^{L}\mathbb{E}\left(\widehat{Y}_{\ell}\right),$
where the correction term $Y_{\ell}$ is defined as in (16). The expectations $\mathbb{E}(\widehat{Y}_{\ell})$ will then be approximated through a Monte Carlo estimator using $N_{\ell}$ samples at the $\ell$-th spatial mesh level. The resulting unbiased estimator $A(\widehat{u}_{L})$ for $\mathbb{E}(u)$ is defined as
$A\left(\widehat{u}_{L}\right):=\sum_{\ell=0}^{L}\frac{1}{N_{\ell}}\sum_{i=1}^{N_{\ell}}\widehat{Y}_{\ell}^{(i)},$ (28)
and satisfies ${\displaystyle\mathbb{E}\left(A(\widehat{u}_{h})\right)=\mathbb{E}(\widehat{u}_{L})}$ and ${\displaystyle\mathbb{V}\left(A(\widehat{u}_{h})\right)=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left(\widehat{Y}_{\ell}^{(i)}\right)}{N_{\ell}}}$. We remark that for the latter property to hold, the samples $\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)$ comprising the term $\widehat{Y}_{\ell}^{(i)}$ for each level $\ell$ are drawn independently of each other (i.e. the samples at any given level are not reused for the subsequent levels). However, within each term $\widehat{Y}_{\ell}^{(i)}$, in view of the identity
$\mathbb{V}\left(\widehat{Y}_{\ell}^{(i)}\right)=\mathbb{V}\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)=\mathbb{V}\left(\widehat{u}_{\ell}^{(i)}\right)+\mathbb{V}\left(\widehat{u}_{\ell-1}^{(i)}\right)-2\text{Cov}\left(\widehat{u}_{\ell}^{(i)},\widehat{u}_{\ell-1}^{(i)}\right)\,,$
the strong correlation between $\widehat{u}_{\ell}^{(i)}$ and $\widehat{u}_{\ell-1}^{(i)}$ as $\ell$ grows leads to a reduction in variance $\mathbb{V}\left(\widehat{Y}_{\ell}^{(i)}\right)$. This contributes to a decrease in the sample size as the mesh gets finer, improving efficiency.
@@
$\mathcal{E}_{A}^{2}$ $=\mathbb{E}\left[\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{L}\right)\right\|^{2}_{Z}\right]+\mathbb{E}\left[\left\|\mathbb{E}\left(\widehat{u}_{L}\right)-A\left(\widehat{u}_{L}\right)\right\|_{Z}^{2}\right]=\underbrace{\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{L}\right)\right\|^{2}_{Z}}_{\mathcal{E}_{\text{Bias}}^{2}}\;+\;\underbrace{\mathbb{V}\left(A\left(\widehat{u}_{L}\right)\right)}_{\mathcal{E}_{\text{Stat}}^{2}}$
$=\left\|\mathbb{E}(u)-\mathbb{E}\left(\widehat{u}_{L}\right)\right\|^{2}_{Z}+\sum_{\ell=0}^{L}\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{N_{\ell}}\leq\underbrace{\bigg{(}\left\|\mathbb{E}(u)-\mathbb{E}(u_{L})\right\|_{Z}+\left\|\mathbb{E}(u_{L})-\mathbb{E}\left(\widehat{u}_{L}\right)\right\|_{Z}\bigg{)}^{2}}_{\left(\mathcal{E}_{\text{Dis}}+\mathcal{E}_{\text{Interp}}\right)^{2}}\;+\;\underbrace{\sum_{\ell=0}^{L}\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{N_{\ell}}}_{\mathcal{E}_{\text{Stat}}^{2}}.$ (29)
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\frac{C_{m}M_{L}^{-\alpha}}{\left\|\mathbb{E}(u_{L})\right\|_{Z}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\frac{C_{p}P^{-\nu}}{\left\|\mathbb{E}(u_{L})\right\|_{Z}}\leq\theta_{2}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Stat}}^{2}=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{N_{\ell}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\leq\theta\epsilon^{2},$ (30)

### v2

@@ 1.1 The Grad-Shafranov free boundary problem with uncertainty
$-\nabla\,\cdot\,\left(\frac{1}{\mu(u,\boldsymbol{\omega})r}\nabla u(\cdot,\boldsymbol{\omega})\right)=\left\{\begin{array}[]{ll}r\frac{d}{du}p(u,\boldsymbol{\omega})+\frac{1}{2\,\mu r}\frac{d}{du}g^{2}(u,\boldsymbol{\omega})&\quad\text{ in }\Omega_{p}(u,\boldsymbol{\omega})\,,\\ I_{k}(\boldsymbol{\omega})/S_{k}&\quad\text{ in }\Omega_{W_{k}}\\ 0&\quad\text{ elsewhere.}\end{array}\right.$ (2)
@@ 1.2 Overview of the article
In its simplest form, a Monte Carlo estimate of the expectation $\mathbb{E}[u]$ gathers a large number, $N$, of the realizations of $u(\cdot,\boldsymbol{\omega})$ and estimates the expectation by the mean of the sample as
$\mathbb{E}[u]\approx\frac{1}{N}\sum_{i=1}^{N}u\left(\cdot,\boldsymbol{\omega}^{(i)}\right)\,.$
Based on the sample mean, the approximation is unbiased and, by the central limit theorem, it will converge to the true value as $N^{-1/2}$. Due to the slow rate of convergence, a large number of samples may need to be gathered to obtain a reasonably accurate approximation of $\mathbb{E}[u]$. Therefore, in simple words, the cost of obtaining a Monte Carlo estimate boils down to the workload associated with sampling. In our case, $u$ maps the stochastic parameters in (2) to the solution of the free boundary problem. In practical terms, the task of obtaining a Monte Carlo approximation for $\mathbb{E}[u]$ entails the use of a computational discretization of the mapping, and the computation of a large number of numerical solutions to (2) for different values of the stochastic parameters. Unfortunately, even if obtaining such a pool of numerical solutions is computationally feasible, the cost of numerically solving the free boundary problem repeatedly can make the process unpractical. The goal then is then to bypass the need for a direct numerical solution of (2) for every new sample gathered, or at least mitigate the work required to obtain each sample.
@@ 2.1 Sparse grid stochastic collocation
where $P$ denotes the number of nodes in the sparse grid, $C$ is a positive constant and the power $\nu$ is an increasing function of the size of the domain of definition of the function’s analytic extension in the complex plane. On the other hand, if the mapping is less regular with respect to the stochastic parameter $\boldsymbol{\omega}$, for instance of class $C^{k}(\boldsymbol{W},Z)$, then [3, Theorem 8]
@@ 2.2 Notational considerations
$\mathbb{V}[u]:=\mathbb{E}\left[\left\|u-\mathbb{E}[u]\right\|_{Z}^{2}\right]\,,$ (8)
@@
where $\alpha$ and $\nu$ are the orders of the sample-wise discretization error and interpolation error in (6). The exponent $\alpha$ depends on both the spatial dimension $d$ and the choice of norm for $Z$. In the two–dimensional case that we are dealing with, the number of spatial grid nodes, $M$, behaves like $M\propto h^{-2}$. Therefore, under the usual (and reasonable) assumption of $H^{2}$–regularity for $u$, the discretization achieves $\mathcal{O}(h^{2})$ convergence in the $L^{2}$ norm and $\alpha=1$ [4, Corollary 7.7].
The constants $C_{m}(\boldsymbol{\omega}^{(i)})$ and $C_{p}(\boldsymbol{\omega}^{(i)})$ depend only on the problem geometry and the particular realization $\boldsymbol{\omega}^{(i)}$. Under the assumption that the solution’s regularity is uniformly controlled across samples, they are both bounded uniformly by deterministic constants $C_{m}$ and $C_{p}$. Specifically, the assumption that the solution’s derivatives satisfy uniform bounds almost surely [1, 2, 32, 37, 38] ensures the finiteness of the bounds
$C_{m}:=\sup_{\boldsymbol{\omega}\in\Omega}C_{m}(\boldsymbol{\omega})\qquad\text{ and }\qquad C_{p}:=\sup_{\boldsymbol{\omega}\in\Omega}C_{p}(\boldsymbol{\omega})\,.$
We remark that, in what follows, when describing the different approaches to surrogate building, the associated values of the constants $C_{m}$, $C_{p}$, $\alpha$, and $\nu$ may vary from section to section but their properties, as enlisted above, remain true.
@@ 2.3 Collocation with a single level of spatial discretization
Considering that, for a nonlinear problem like the one in our application, the computational cost of obtaining a sample is itself a random variable, we will estimate the computational work required to build this kind of surrogate, by assuming that the average work per sample is of the form $W\simeq M^{\gamma}$ for some $\gamma>0$ that depends on the particular method used to solve the system of equations arising from the discretization, and $A\simeq B$ means $C_{1}B\leq A\leq C_{2}B$ for positive $A$ and $B$, with constants $C_{1},C_{2}$ independent of sample size $N$ and number of spatial grid nodes $M$. Hence, making use of (12), for a surrogate built using $P$ sparse grid nodes, the expected work to construct the surrogate, measured in terms of $\epsilon$, is given by
@@
where the subscript “SL-SL” denotes the surrogate built with a single-level sparse grid and a single-level spatial grid, the superscript “off” represents an offline process, and, to recall, $\epsilon$ represents the desired accuracy, $\alpha$ is the convergence rate of the spatial discretization, $\nu$ is the convergence rate of the interpolant, $\gamma$ is the rate associated with the non-linear solver that characterizes the average work per sample.
@@ 2.4 Collocation with multilevel spatial discretization
$\left\|Y_{\ell}^{(i)}-\widehat{Y}_{\ell}^{(i)}\right\|_{Z}=\left\|\left(u_{\ell}^{(i)}-u_{\ell-1}^{(i)}\right)-\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)\right\|_{Z}=\left\|\left(u_{\ell}^{(i)}-\widehat{u}_{\ell}^{(i)}\right)-\left(u_{\ell-1}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)\right\|_{Z}$
@@
These intuitive observations in fact hold for the case of Lagrangian interpolation and Smolyak grids based on nested quadrature points—which include the Clenshaw-Curtis strategy we employ in this paper. Moreover, it was shown in [37, Section 5] that there are in fact positive constants $C^{(i)},\rho$, and $\nu$ such that
@@ 3.1 Sampling with a single-spatial level surrogate
$A\left[\widehat{u}_{h}\right]:=\frac{1}{N}\sum_{i=1}^{N}\widehat{u}_{h}^{(i)}.$
@@
$\mathbb{E}[A[\widehat{u}_{h}]]=\mathbb{E}[\widehat{u}_{h}]\qquad\text{ and }\qquad\mathbb{V}[A[\widehat{u}_{h}]]=\mathbb{V}[\widehat{u}_{h}]/N\,,$ (23)
where $\mathbb{V}[\cdot]$ is defined as in (8). To assess the accuracy of the estimator we will use the mean squared error (MSE) defined as
$\mathcal{E}_{A}^{2}:=\mathbb{E}\left[\left\|\mathbb{E}[u]-A\left[\widehat{u}_{h}\right]\right\|_{Z}^{2}\right]\,$
@@
$\mathcal{E}_{A}^{2}=\mathbb{E}\left[\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}\right]+\mathbb{E}\left[\left\|\mathbb{E}\left[\widehat{u}_{h}\right]-A\left[\widehat{u}_{h}\right]\right\|_{Z}^{2}\right]=\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}+\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N}=\mathcal{E}_{\text{Bias}}^{2}+\mathcal{E}_{\text{Stat}}^{2},$
@@
$\mathcal{E}_{\text{Bias}}^{2}:=\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}\qquad\text{ and }\qquad\mathcal{E}_{\text{Stat}}^{2}:=\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N}.$
The bias term can be further bounded by the sum of discretization error, $\mathcal{E}_{\text{Dis}}:=\|\mathbb{E}[u]-\mathbb{E}[u_{h}]\|_{Z}$, and interpolation error  $\mathcal{E}_{\text{Interp}}:=\left\|\mathbb{E}[u_{h}]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|_{Z}$, as
$\mathcal{E}_{\text{Bias}}=\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|_{Z}\leq\|\mathbb{E}[u]-\mathbb{E}[u_{h}]\|_{Z}+\left\|\mathbb{E}[u_{h}]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|_{Z}=\mathcal{E}_{\text{Dis}}+\mathcal{E}_{\text{Interp}}\,.$
@@
If the expectation of $u$ is nonzero, the normalized mean squared error (nMSE), denoted ¯$\overline{\mathcal{E}}_{A}^{2}$, is considered. A similar normalization and notation is used for the discretization, interpolation and statistical errors. When the normalizing factor $\left\|\mathbb{E}[u]\right\|_{Z}^{2}$ is not available or computable we approximate it by $\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}$. For instance the nMSE is approximated as
¯$\overline{\mathcal{E}}_{A}^{2}\approx\frac{\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}}{\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}+\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}\leq\left(\overline{\mathcal{E}}_{\text{Dis}}+\overline{\mathcal{E}}_{\text{Interp}}\right)^{2}+\overline{\mathcal{E}}_{\text{Stat}}^{2}.$
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\theta_{2}\epsilon,\qquad\text{ and }\qquad\overline{\mathcal{E}}_{\text{Stat}}=\sqrt{\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}}\leq\sqrt{\theta}\epsilon,$ (24)
where $\theta\in(0,1)$ is the splitting parameter between the relative bias and the statistical errors, and $\theta_{1}$ and $\theta_{2}$ are the splitting parameters between the discretization and interpolation errors, such that $\theta_{2}=\sqrt{1-\theta}-\theta_{1}$. The number of grid points in the physical space $M$ and the parameter space $P$ are determined by (12) with a scaling factor $\left\|\mathbb{E}[u_{h}]\right\|_{Z}$, while the sample size $N$ is estimated from (24) as
$N=\left\lceil\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}\right\rceil\simeq\epsilon^{-2}.$ (25)
@@ 3.2 Sampling with surrogates with multilevel spatial discretization
We express the surrogate at the finest spatial grid level $L$ as a telescoping sum of surrogates on coarser grids, and then exploit the linearity of the expectation to approximate $\mathbb{E}[u_{h}]$ as
$\mathbb{E}[u_{h}]\approx\mathbb{E}\left[\widehat{u}_{L}\right]=\sum_{\ell=0}^{L}\mathbb{E}[\widehat{u}_{\ell}-\widehat{u}_{\ell-1}]=\sum_{\ell=0}^{L}\mathbb{E}\left[\widehat{Y}_{\ell}\right],$
where the correction term $Y_{\ell}$ is defined as in (16). The expectations $\mathbb{E}[\widehat{Y}_{\ell}]$ will then be approximated through a Monte Carlo estimator using $N_{\ell}$ samples at the $\ell$-th spatial mesh level. The resulting unbiased estimator $A[\widehat{u}_{L}]$ for $\mathbb{E}[u]$ is defined as
$A\left[\widehat{u}_{L}\right]:=\sum_{\ell=0}^{L}\frac{1}{N_{\ell}}\sum_{i=1}^{N_{\ell}}\widehat{Y}_{\ell}^{(i)},$ (28)
and satisfies ${\displaystyle\mathbb{E}\left[A[\widehat{u}_{h}]\right]=\mathbb{E}[\widehat{u}_{L}]}$ and ${\displaystyle\mathbb{V}\left[A[\widehat{u}_{h}]\right]=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}^{(i)}\right]}{N_{\ell}}}$. We remark that for the latter property to hold, the samples $\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)$ comprising the term $\widehat{Y}_{\ell}^{(i)}$ for each level $\ell$ are drawn independently of each other (i.e. the samples at any given level are not reused for the subsequent levels). However, within each term $\widehat{Y}_{\ell}^{(i)}$, in view of the identity
$\mathbb{V}\left[\widehat{Y}_{\ell}^{(i)}\right]=\mathbb{V}\left[\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right]=\mathbb{V}\left[\widehat{u}_{\ell}^{(i)}\right]+\mathbb{V}\left[\widehat{u}_{\ell-1}^{(i)}\right]-2\text{Cov}\left[\widehat{u}_{\ell}^{(i)},\widehat{u}_{\ell-1}^{(i)}\right]\,,$
the strong correlation between $\widehat{u}_{\ell}^{(i)}$ and $\widehat{u}_{\ell-1}^{(i)}$ as $\ell$ grows leads to a reduction in variance $\mathbb{V}\left[\widehat{Y}_{\ell}^{(i)}\right]$. This contributes to a decrease in the sample size as the mesh gets finer, improving efficiency.
@@
$\mathcal{E}_{A}^{2}$ $=\mathbb{E}\left[\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{L}\right]\right\|^{2}_{Z}\right]+\mathbb{E}\left[\left\|\mathbb{E}\left[\widehat{u}_{L}\right]-A\left[\widehat{u}_{L}\right]\right\|_{Z}^{2}\right]=\underbrace{\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{L}\right]\right\|^{2}_{Z}}_{\mathcal{E}_{\text{Bias}}^{2}}\;+\;\underbrace{\mathbb{V}\left[A\left[\widehat{u}_{L}\right]\right]}_{\mathcal{E}_{\text{Stat}}^{2}}$
$=\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{L}\right]\right\|^{2}_{Z}+\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{N_{\ell}}\leq\underbrace{\bigg{(}\left\|\mathbb{E}[u]-\mathbb{E}[u_{L}]\right\|_{Z}+\left\|\mathbb{E}[u_{L}]-\mathbb{E}\left[\widehat{u}_{L}\right]\right\|_{Z}\bigg{)}^{2}}_{\left(\mathcal{E}_{\text{Dis}}+\mathcal{E}_{\text{Interp}}\right)^{2}}\;+\;\underbrace{\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{N_{\ell}}}_{\mathcal{E}_{\text{Stat}}^{2}}.$ (29)
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\frac{C_{m}M_{L}^{-\alpha}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\frac{C_{p}P^{-\nu}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{2}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Stat}}^{2}=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{N_{\ell}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\leq\theta\epsilon^{2},$ (30)